OpenStudy (unofficialllyy):

An astronaut, including her spacesuit, has a mass of 150kg at a NASA training facility on Earth. She then travels to Mars, where the acceleration due to gravity is 3.3 m/s/s. What is her mass on Mars? What is her weight on Mars?

OpenStudy (unofficialllyy):

@brainlesslonelygirl

OpenStudy (brainlesslonelygirl):

convert the mass from kg to g first thing = 150*1000= 150000g now, mass is the amount of matter in a body and it remains the same no matter where you are. It is unaffected by the force of gravity. so the mass on mars is the same as the mass on earth. the equation for weight is W=mg where g is the gravitational force which varies planet to planet plugging the quantities in the equation we get W= 150000*3.3= 495000 which you can again convert if you want to which will give you W= 495kg

OpenStudy (unofficialllyy):

I tried that and it's wrong.

OpenStudy (brainlesslonelygirl):

welp.

OpenStudy (unofficialllyy):

@mathmale